OpenStudy (anonymous):

How to find the point of f(x)= (3x-2x^2)^2 where the tangent line is parallel to the x-axis?

OpenStudy (anonymous):

firtst of all you need to find then slope of the tangent by dfferentiating f(x), do you know how to do that?

OpenStudy (anonymous):

or what you could do is factorize f(x)

OpenStudy (anonymous):

for a line to be parallel to the x axis then x=0

OpenStudy (anonymous):

so plug zeros in the x's of f(x)

OpenStudy (anonymous):

y=? that would be parallel to the x-axis and this is the third method you can use...

OpenStudy (anonymous):

i used to product rule and got f' (x) to be 16x^3 -12x^2-6x. I don't know if its right thought, because the answers are (0,0) (3/2, 0) and (3/4, 81/64)

OpenStudy (anonymous):

no thats wrong

OpenStudy (anonymous):

ok. what would it be though?

OpenStudy (anonymous):

\[f(x)= (3x + -2x^2)^2\]

OpenStudy (anonymous):

the best way with less error to solve this function is to expand it...

OpenStudy (anonymous):

then differentiate it

OpenStudy (anonymous):

\[(a+b)^2 = a^2 +2ab+b^2\]

OpenStudy (anonymous):

(3x-2x^2)(3x-2x^2). But when I used the product rule to differentiate it, i got the above equation

OpenStudy (anonymous):

well i would do it another way \[f(x)= (3x + -2x^2)^2 = 9x^2-6x^3+4x^4\]

OpenStudy (anonymous):

now you can differentiate it simply

OpenStudy (anonymous):

then would i use the power rule to differentiate it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i would get 16x^3-18x^2+18x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but i don't get how to get the other two points as the answer. I get that if i sub 0 in, then i will get 0. but I don't get how to get (3/2, 0) and (3/4, 81/64)

OpenStudy (anonymous):

you need to factor the third order polynomial...

OpenStudy (anonymous):

can I long divide it, if i find an x value that makes it equal to 0

OpenStudy (anonymous):

16x^3-18x^2+18x= 2x(8x^2-9x+9)

OpenStudy (anonymous):

i'm sorry, but i still don't know how to get those values

OpenStudy (anonymous):

well you use the quaratic equation

OpenStudy (anonymous):

from what i can see you are going to get a complex number, have you done that before...

OpenStudy (anonymous):

i tried, i get a negative under the square root

OpenStudy (anonymous):

yes that is a complex number...

OpenStudy (anonymous):

ok do you know the chain rule

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then you need to check if the function is correct

OpenStudy (anonymous):

ok . I guess i will just ask my teacher

OpenStudy (anonymous):

ok find the 2nd derivative

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i didn't learn that yet

OpenStudy (anonymous):

ok there is a simpler way

OpenStudy (anonymous):

i will just ask my teacher, it ok. thanks

OpenStudy (anonymous):

well you know that x=0, right the one we factored out

OpenStudy (anonymous):

so we said when x =0 that will give you the value of y, and we get y=0

OpenStudy (anonymous):

so the line passing through the piont(0,0) is the line that is parallel to the x-axis

OpenStudy (anonymous):

based on basic geomatry thoerem

OpenStudy (anonymous):

i don't know what that is

OpenStudy (anonymous):

as i said at the beginig that when x=0 y=?

OpenStudy (anonymous):

y = 0

OpenStudy (anonymous):

ok let me put it this way set y =0, what values for x would you get?

OpenStudy (anonymous):

(3x-2x^2)^2 = 0 -> x (3-2x)=0

OpenStudy (anonymous):

what is x =?

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

so we get the point (3/2, 0)

OpenStudy (anonymous):

did u get that?

OpenStudy (anonymous):

\[\frac{y-y_1}{x-x_1}=m\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now remeber a line parallel to an axis tells you that the slop is =0 because there is no slope right, its a straight line...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know how to complete the square

OpenStudy (anonymous):

because when you complete the square you'll find the vertix which would give you the point parallel to the x axis

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok good find the square and i will see you later because its too late here 5am in the morning ...

OpenStudy (anonymous):

take care and good lcuk

OpenStudy (anonymous):

ok, thank you